# Measure-Theoretic Time-Delay Embedding

### v1

@@ Abstract
The celebrated Takens’ embedding theorem provides a theoretical foundation for reconstructing the full state of a dynamical system from partial observations. However, the classical theorem assumes that the underlying system is deterministic and that observations are noise-free, limiting its applicability in real-world scenarios. Motivated by these limitations, we rigorously establish a measure-theoretic generalization that adopts an Eulerian description of the dynamics and recasts the embedding as a pushforward map between probability spaces. Our mathematical results leverage recent advances in optimal transportation theory. Building on our novel measure-theoretic time-delay embedding theory, we have developed a new computational framework that forecasts the full state of a dynamical system from time-lagged partial observations, engineered with better robustness to handle sparse and noisy data. We showcase the efficacy and versatility of our approach through several numerical examples, ranging from the classic Lorenz-63 system to large-scale, real-world applications such as NOAA sea surface temperature forecasting and ERA5 wind field reconstruction.
@@ 1 Introduction
In situations where the full state is not directly observable, time-delay embedding has become a fundamental technique in the analysis of dynamical systems. This method involves concatenating time-lagged versions of a scalar time-dependent observation into a high-dimensional state vector, resulting in a system that is topologically equivalent to the full, unobserved dynamics. The celebrated Takens’ embedding theorem supplies the theoretical foundation for time-delay embedding and has inspired a myriad of works over the last four decades that perform data-driven analysis on partially observed nonlinear systems [50, 27]. Notable applications of time-delay embedding include forecasting [4, 58], noise reduction [22, 29], control [37], as well as various biological studies [57, 40, 46].
@@
Given these practical challenges, it is necessary to consider variants of time-delay embedding that incorporate assumptions of randomness. Notably, it was shown in [44] that the time-delay map is an embedding almost surely in the space of observation functions, and in [5] that the delay map formed by a polynomially perturbed observation function is an embedding at almost all points. However, these works do not address intrinsic or extrinsic noise in the state itself. On the other hand, [14] leverages statistical techniques to quantify the effect of i.i.d. extrinsic noise on state reconstruction and time-series prediction. Approaches incorporating stochasticity often face the curse of dimensionality. For example, [47, 48] showed that Takens’ theorem holds for stochastically forced systems whose vector fields lie in finite discrete probability spaces. Nevertheless, the embedding dimension increases with the dimension of the probability space, and for general stochastic systems with an infinite-dimensional sample space, such variants of time-delay embedding no longer hold.
@@
Our second main contribution is leveraging the measure-theoretic time-delay embedding theory to establish a robust computational framework for learning the inverse embedding function ($\Phi^{-1}$) from data. This function, known as the full-state reconstruction map, typically has no analytical form but is crucial for forecasting the complete state of a nonlinear system from partially observed data. To learn the reconstruction map, we first extract empirical measures in both the full-state space ($M$) and the delay space ($N$) based on a single time trajectory. We then select a suitable metric over the probability space as the objective function and enforce that the pushforwards of the empirical measures in the delay space match the corresponding measures in the full-state space. In contrast with our approach, existing methods commonly learn the reconstruction map using the pointwise mean-squared error loss [4, 40, 58]. When the observed data is sparse and noisy, the accuracy of such pointwise methods can be compromised, while our measure-theoretic approach demonstrates clear robustness. In Section 4, we showcase the effectiveness of our proposed approach through various numerical tests on classic synthetic examples, such as the Lorenz-63 and Lotka–Volterra systems, as well as large-scale real-world applications, including the reconstruction of NOAA Sea Surface Temperature [41] and ERA5 wind speed datasets from partial observations [24].
@@ 2.1 Takens’ Embedding Theorem
    If $v(x)=0$, then the eigenvalues of $(d\phi_{\tau})_{x}:T_{x}M\rightarrow T_{\phi_{\tau}(x)}M$ are all different, and none of them equals $1$.
@@
Table 1: A list of mathematical notations in Section 2 and Section 3.
In (1), $m\in\mathbb{N}$ is known as the embedding dimension, and $\tau>0$ is the so-called time-delay. By “generic” we mean that the collection of observation functions $h\in C^{2}(M,\mathbb{R})$ for which (1) defines an embedding is open and dense in the $C^{2}$ topology. In particular, when (1) is an embedding, the delayed trajectory $\{\Phi_{h,\phi_{\tau}}(\phi_{t}(x))\}_{t\geq 0}$ is topologically equivalent to the original orbit $\{\phi_{t}(x)\}_{t\geq 0}\subseteq M$. This perspective has motivated the use of delay coordinates in various applications, such as time-series prediction [17], attractor reconstruction [38], causality detection [49], and noise reduction [22, 29].
@@ 2.2 Optimal Transport and the Wasserstein Geometry
Definition 2 (The pushforward operator [1]).
@@ 2.2.1 The continuity equation
Definition 3 (Absolute continuity of curves and maps [1]).
@@
Definition 4 (Metric derivative [1]).
@@
exists $L^{1}$ a.e. in $t$, and is called the metric derivative of $\rho_{t}$.
In Theorem 1, the flow $\phi_{t}$ is generated by a smooth vector field $v:M\rightarrow TM$, such that at every point $\frac{d}{dt}\phi_{t}(x)=v(x)$. In contrast, in the Wasserstein space, the curve $\rho_{t}$ is generated by a square-integrable vector field
$v_{t}$ $\in L^{2}(TM,\rho_{t})=\Big{\{}v:[0,T]\times M\rightarrow TM,\,$
@@
where $g_{x}$ is the Riemannian metric on $M$. Note that $v_{t}(x)\in T_{x}M$ and $g_{x}:T_{x}M\times T_{x}M\rightarrow\mathbb{R}$. This vector field $v_{t}$ has to additionally satisfy the continuity equation ((4)) in the distributional sense (see (5)):
$\frac{\partial\rho_{t}}{\partial t}+\nabla\cdot(v_{t}\rho_{t})=0\,.$ (4)
To be more precise,
@@
for all $\varphi\in\mathcal{D}(M\times[0,T])\,$ where $\mathcal{D}(M\times[0,T])$ denotes the space of test functions on $M\times[0,T]$ (see Remark 2). This continuity equation comes from the conservation of mass. In the rest of the paper, when we say that the continuity equation is satisfied, we mean (5) holds. Theorem 2 below shows that any AC curve is a solution to (5) for an appropriate choice of $v_{t}$ [1, Chapter 8].
@@
Let $v_{t}$ be a Borel vector field such that:
@@
    $\int_{0}^{T}\sup\limits_{B}\left(|v_{t}|+\text{Lip}(v_{t},B)\right)\,{\mathrm{d}}t<\infty$ ,
where $\text{Lip}(v_{t},B)$ denotes the Lipscitz constant of $v_{t}$ on the set $B\in M$. Let $\phi_{t}:M\times[0,T]\rightarrow M$ be the solution to the ODE
$\phi_{t}(x,s)=x,\qquad\frac{d}{dt}\phi_{t}(x,s)=v_{t}(\phi_{t}(x,s)).$ (6)
Suppose that for $\rho_{0}$-a.e. $x\in M$, $t<\sup(I(x,0))$ for $t\in[0,T]$, where $I(x,0)$ denotes the interval on which solutions to (6) at $s=0$ and position $x$ exist. Then,
$\rho_{t}=\phi_{t}\#\rho_{0}$
solves the continuity equation. Conversely, let $v_{t}$ be a Borel vector field satisfying conditions (i) and (ii). Let $\rho\in\mathcal{P}_{2}(M)$ and define $\rho_{t}=\phi_{t}\#\rho$, where $\phi_{t}$ is the flow of the system $\dot{x}=v_{t}(x)$. Then $\rho_{t}$ is the unique solution to the continuity equation with initial condition $\rho$.
@@
Without loss of generality, we will work with test functions that are compactly supported and continuously differentiable on $M\times[0,T]$, i.e., $\mathcal{D}(M\times[0,T])=\mathcal{C}^{1}_{c}(M\times[0,T])$. This gives us sufficient regularity to define all the distributional derivatives of $\rho$ and $v_{t}$. Moreover, it can be shown (see [1, Remark 8.1.1]) that if (5) holds for any $\varphi\in\mathcal{D}(M\times[0,T])$, it also holds for $\varphi\in\mathcal{C}^{\infty}_{c}(M\times[0,T])$. Similarly, for time-independent test functions, we can choose $\mathcal{D}(M)=\mathcal{C}^{1}_{c}(M)$. The same extends to test functions for $N$ and $N\times[0,T]$.
@@ 2.2.2 The tangent space of $\mathcal{P}_{2}(M)$
Hereafter, we assume $(M,g)$ is a Riemannian manifold and $g$ is the Riemannian metric. The tangent space to a manifold at a specific point is comprised of all possible infinitesimal directions of motion starting from that point. The continuity equation ((4)) tells us how a measure evolves in time along the direction specified by a vector field $v_{t}$, and Theorem 2 verifies the existence of such directions. Hence, intuitively, we may define $T_{\rho}\mathcal{P}_{2}(M)=L^{2}(TM,\rho)$ as the tangent space to $\mathcal{P}_{2}(M)$ at $\rho$. However, not every element in $L^{2}(TM,\mu)$ generates a different flow in $\mathcal{P}_{2}(M)$.
@@
Fix an AC curve $\rho_{t}$ and consider $v_{t},w_{t}\in L^{2}(TM,\rho_{t})$ where $\nabla\cdot(w_{t}\rho_{t})=0$ and $v_{t}$ satisfies $\partial_{t}\rho_{t}+\nabla\cdot(v_{t}\rho_{t})=0$ (the existence of $v_{t}$ is guaranteed by Theorem 2). Then one can check that $\rho_{t}$ also satisfies $\partial_{t}\rho_{t}+\nabla\cdot((v_{t}+w_{t})\rho_{t})=0$. Hence, any two vector fields that differ by a $\rho_{t}$-weighted divergence-free component produce the same solution to the continuity equation. We say a vector field $v$ is $\rho$-weighted divergence-free if $\nabla\cdot(v\rho)=0$ holds in the distributional sense, i.e., $\forall\varphi\in\mathcal{D}(M):\int_{M}g_{x}\left(\nabla\varphi(x),v(x)\right)\,{\mathrm{d}}\rho(x)=0$. A time-dependent vector field $v_{t}$ is divergence free if the above holds for $L^{1}$-a.e. $t$.
@@ 2.2.3 The projection operator
where $T_{\rho}^{\perp}\mathcal{P}_{2}(M)$ is the orthogonal complement of $T_{\rho}\mathcal{P}_{2}(M)$ with respect to the $\rho$-weighted $L^{2}$ inner product, and $P^{\rho}:L^{2}(TM,\rho)\rightarrow T_{\rho}\mathcal{P}_{2}(M)$ is the projection operator. We will use this projection operator to obtain the unique vector field defining the derivative of a map between probability spaces. The kernel of $P^{\rho}$ comprises all $\rho$-weighted divergence-free vector fields $w$ such that $\nabla\cdot(\rho w)=0$ in the distributional sense. One important property of $P^{\rho}$ is given in Proposition 1, whose complete proof is in the supplementary materials.
@@
2.2.4 Metric change of variables formula
We review the following result from Riemannian geometry, which will be used to show the injectivity of the metric derivative operator (see the proof of Theorem 4). A complete proof appears in the supplementary materials.
@@ 2.2.4 Metric change of variables formula
$g_{x}(\nabla(\varphi\circ f)(x),X(x))=q_{f(x)}\Big{(}(\nabla\varphi)(f(x)),df_{x}X(x)\Big{)}.$ (9)
@@ 3.1 Differentiable curves in $\mathcal{P}_{2}(M)$
On the Lagrangian level, given an initial condition $x_{0}$, the flow $\phi_{t}(x_{0})$ generates a differentiable curve $x(t)=\phi_{t}(x_{0})$ in $M$. Similarly, on the Eulerian level, we want to have a differentiable curve $\rho_{t}$ and a vector field $v_{t}\in T_{\rho_{t}}\mathcal{P}_{2}(M)$ such that they satisfy the continuity equation ((4)). To begin with, we define the notion of a vector field along a curve in $\mathcal{P}_{2}(M)$.
@@
Consider a curve $\rho_{t}\in\mathcal{P}_{2}(M)$. We say that $v_{t}:[0,1]\rightarrow T_{\rho_{t}}\mathcal{P}_{2}(M)$ is a vector field along the curve $\rho_{t}$ if the tuple $(\rho_{t},v_{t})$ satisfies the continuity equation ((4)). If such a vector field exists, we denote it by $v_{t}:=\frac{d}{dt}\rho_{t}$.
@@
The proof appears in the supplemental materials and relies on the following Proposition:

### v2

@@ Abstract
The celebrated Takens’ embedding theorem provides a theoretical foundation for reconstructing the full state of a dynamical system from partial observations. However, the classical theorem assumes that the underlying system is deterministic and that observations are noise-free, limiting its applicability in real-world scenarios. Motivated by these limitations, we formulate a measure-theoretic generalization that adopts an Eulerian description of the dynamics and recasts the embedding as a pushforward map between spaces of probability measures. Our mathematical results leverage recent advances in optimal transport. Building on the proposed measure-theoretic time-delay embedding theory, we develop a computational procedure that aims to reconstruct the full state of a dynamical system from time-lagged partial observations, engineered with robustness to handle sparse and noisy data. We evaluate our measure-based approach across several numerical examples, ranging from the classic Lorenz-63 system to real-world applications such as NOAA sea surface temperature reconstruction and ERA5 wind field reconstruction.
keywords:
dynamical systems, time-delay embedding, optimal transport , state reconstruction
@@ 1 Introduction
In situations where the full state is not directly observable, time-delay embedding has become a fundamental technique in the analysis of dynamical systems. This method involves concatenating time-lagged versions of a scalar time-dependent observation into a high-dimensional state vector, resulting in a system that is topologically equivalent to the full, unobserved dynamics. The celebrated Takens’ embedding theorem supplies the theoretical foundation for time-delay embedding and has inspired a myriad of works over the last four decades that perform data-driven analysis on partially observed nonlinear systems [8, 9]. Notable applications of time-delay embedding include forecasting [10, 11], noise reduction [12, 13], control [14], as well as various biological studies [4, 15, 16]. The theory of time-delay embeddings has been extended beyond classical reconstruction to address problems of prediction and modeling of dynamical evolution [17]. It has been shown that reliable short-term forecasts can be achieved even when embeddings contain self-intersections [18], while subsequent work established a rigorous mathematical framework for prediction from such non-injective embeddings [19]. More recently, practical tests for assessing topological conjugacy directly from time series have been proposed [20].
@@
Given these practical challenges, it is necessary to consider variants of time-delay embedding that incorporate assumptions of randomness. Notably, it was shown in [21] that the time-delay map is an embedding almost surely in the space of observation functions, and in [22] that the delay map formed by a polynomially perturbed observation function is an embedding at almost all points. However, these works do not address intrinsic or extrinsic noise in the state itself. On the other hand, [23] leveraged statistical techniques to quantify the effect of i.i.d. extrinsic noise on state reconstruction and time-series prediction. Moreover, the works [24, 25] focus on the embedding properties of individual Lagrangian trajectories subject to randomness and forcing.
@@
Our second main contribution is leveraging the measure-theoretic time-delay embedding theory to establish a robust computational framework for learning the inverse embedding function ($\Phi^{-1}$) from data. This function, known as the full-state reconstruction map, typically has no analytical form but is crucial for reconstructing the complete state of a nonlinear system from partially observed data. To learn the reconstruction map, we first extract empirical measures in both the full-state space ($M$) and the delay space ($N$) based on a single time trajectory. We then select a suitable metric over the probability space as the objective function and enforce that the pushforwards of the empirical measures in the delay space match the corresponding measures in the full-state space. In contrast with our approach, existing methods commonly learn the reconstruction map using the pointwise mean-squared error loss [10, 15, 11]. When the observed data is sparse and noisy, the accuracy of such pointwise methods can be compromised, while our measure-theoretic approach demonstrates better robustness. In Section 4, we showcase the effectiveness of our proposed approach through various numerical tests on classic synthetic examples, such as the Lorenz-63 and Lotka–Volterra systems, as well as large-scale real-world applications, including the reconstruction of NOAA Sea Surface Temperature [27] and ERA5 wind speed datasets from partial observations [28].
There have been recent advances in leveraging delay-coordinate liftings to apply linear tools to nonlinear dynamics, including kernel integral operators that concentrate the Koopman point spectrum, the forced-linear HAVOK decomposition, and a Koopman–Takens theorem for linear least-squares prediction in the infinite-delay limit [29, 30, 31]. In parallel, measure-theoretic approaches employ optimal transport to construct transfer operators and develop reduced-order models via displacement interpolation and augmentation [32, 33]. Our framework links these efforts by lifting delay embedding at the functional level: we pass from trajectories to probability measures, model evolution as pushforwards on the probability space, and obtain a smooth, one-to-one distributional embedding. This Eulerian lifting complements Koopman/transfer-operator analyses and yields stable reconstruction and forecasting frameworks under finite, noisy, and sparsely sampled time-lagged observations.
@@ 2.1 Takens’ Embedding Theorem
    If $v(x)=0$, then the eigenvalues of $(d\phi_{\tau})_{x}:T_{x}M\rightarrow T_{x}M$ are all different, and none of them equals $1$.
@@
Table 1: A list of mathematical notations in Sections 2 and 3.
In (1), $m\in\mathbb{N}$ is known as the embedding dimension, and $\tau>0$ is the so-called time-delay. By “generic” we mean that the collection of observation functions $h\in\mathcal{C}^{2}(M,\mathbb{R})$ for which (1) defines an embedding is open and dense in the $\mathcal{C}^{2}$ topology. In particular, when (1) is an embedding, the delayed trajectory $\{\Phi_{h,\phi_{\tau}}(\phi_{t}(x))\}_{t\geq 0}$ is topologically equivalent to the original orbit $\{\phi_{t}(x)\}_{t\geq 0}\subseteq M$. This perspective has motivated the use of delay coordinates in various applications, such as time-series prediction [36], attractor reconstruction [37], causality detection [38], and noise reduction [12, 13].
@@ 2.2 Optimal Transport and the Wasserstein Geometry
Definition 2 (The pushforward operator [44, Sec. 5.2]).
@@ 2.2.1 The Continuity Equation
Definition 3 (Absolute continuity of curves and maps [44, Definition 1.1.1]).
@@
Definition 4 (Metric derivative [44, Theorem 1.1.2]).
@@
exists a.e. in $t$, and the map $t\mapsto|\rho_{t}^{\prime}|$ is $L^{1}$. We call $|\rho_{t}^{\prime}|$ the metric derivative of $\rho_{t}$.
In Theorem 1, the flow $\phi_{t}$ is generated by a smooth vector field $v:M\rightarrow TM$, such that at every point, $\frac{d}{dt}\phi_{t}(x)=v(x)$. In contrast, in the Wasserstein space, the curve $\rho_{t}$ is generated by a square-integrable vector field
$v_{t}$ $\in L^{2}(TM,\rho_{t})=\Big\{v_{t}:M\rightarrow TM,\,t\in[0,T],\,$
@@
where $g_{x}$ is the Riemannian metric on $M$. Note that $v_{t}(x)\in T_{x}M$ and $g_{x}:T_{x}M\times T_{x}M\rightarrow\mathbb{R}$. This vector field $v_{t}$ has to additionally satisfy the continuity equation
$\frac{\partial\rho_{t}}{\partial t}+\nabla\cdot(v_{t}\rho_{t})=0\,$, (4)
in the distributional sense, i.e.,
@@
for all $\varphi\in\mathcal{D}(M\times[0,T])\,$ where $\mathcal{D}(M\times[0,T])$ denotes the space of test functions on $M\times[0,T]$ (see Remark 2). This continuity equation comes from the conservation of mass. In the rest of the paper, when we say that the continuity equation is satisfied, we mean (5) holds. Theorem 2 below shows that any AC curve is a solution to (5) for an appropriate choice of $v_{t}$ [44, Chapter 8].
@@
Let $v_{t}\in L^{2}(TM,\mu_{t})$ be a Borel vector field such that:
@@
    $\int_{0}^{T}\sup\limits_{{B\subset M}}\left(|v_{t}|+\text{Lip}(v_{t},B)\right)\,{\mathrm{d}}t<\infty$ ,
where $\text{Lip}(v_{t},B)$ denotes the Lipschitz constant of $v_{t}$ on the set $B\subset M$. Let $\phi:M\times[0,T]\rightarrow M$ be the solution to the ODE
$\phi(x,0)=x,\qquad\frac{d}{dt}\phi(x,t)=v_{t}\left(\phi(x,t)\right).$ (6)
We will denote $\phi_{t}(x):=\phi(x,t)$. Let $I(x)\subset\mathbb{R}$ be the time interval on which solutions to (6) starting with initial position $x$ exist. Suppose that $0,T]\subseteq I(x)$. Then,
  1. a.
    (Existence) $\rho_{t}=\phi_{t}\#\rho_{0}$ solves the continuity equation (4) with initial condition $\rho_{0}$.
  2. b.
    (Uniqueness) Given any $\rho\in\mathcal{P}_{2}(M)$, $\rho_{t}=\phi_{t}\#\rho$ is the unique solution to the continuity equation with the initial condition $\rho$.
@@
Without loss of generality, we will work with test functions that are compactly supported and continuously differentiable on $M\times[0,T]$, i.e., $\mathcal{D}(M\times[0,T])=\mathcal{C}^{1}_{c}(M\times[0,T])$. This gives us sufficient regularity to define all the distributional derivatives of $\rho$ and $v_{t}$ (see [44, Remark 8.1.1]). No generality is lost because, if (5) holds for any $\varphi\in\mathcal{D}(M\times[0,T])$, it also holds for $\varphi\in\mathcal{C}^{\infty}_{c}(M\times[0,T])$. Similarly, for time-independent test functions, we can choose $\mathcal{D}(M)=\mathcal{C}^{1}_{c}(M)$. The same extends to test functions for $N$ and $N\times[0,T]$.
@@ 2.2.2 The Tangent Space of $\mathcal{P}_{2}(M)$
Hereafter, we assume $(M,g)$ is a Riemannian manifold and $g$ is the Riemannian metric. The tangent space to a manifold at a specific point is comprised of all possible infinitesimal directions of motion starting from that point. The continuity equation (4) tells us how a measure evolves in time along the direction specified by a vector field $v_{t}$, and Theorem 2 verifies the existence of such directions. Hence, intuitively, we may define $T_{\rho}\mathcal{P}_{2}(M)=L^{2}(TM,\rho)$ as the tangent space to $\mathcal{P}_{2}(M)$ at $\rho$. However, not every element in $L^{2}(TM,\mu)$ generates a different flow in $\mathcal{P}_{2}(M)$.
@@
Fix an AC curve $\rho_{t}$ and consider $v_{t},w_{t}\in L^{2}(TM,\rho_{t})$ where $\nabla\cdot(w_{t}\rho_{t})=0$ and $v_{t}$ satisfies $\partial_{t}\rho_{t}+\nabla\cdot(v_{t}\rho_{t})=0$ (the existence of $v_{t}$ is guaranteed by Theorem 2). Then one can check that $\rho_{t}$ also satisfies $\partial_{t}\rho_{t}+\nabla\cdot((v_{t}+w_{t})\rho_{t})=0$. Hence, any two vector fields that differ by a $\rho_{t}$-weighted divergence-free component produce the same solution to the continuity equation. We say a vector field $v$ is $\rho$-weighted divergence-free if $\nabla\cdot(v\rho)=0$ holds in the distributional sense, i.e., $\forall\varphi\in\mathcal{D}(M):\int_{M}g_{x}\left(\nabla\varphi(x),v(x)\right)\,{\mathrm{d}}\rho(x)=0$. A time-dependent vector field $v_{t}$ is divergence free if the above holds for a.e. $t$.
@@ 2.2.3 The Projection Operator
where $T_{\rho}^{\perp}\mathcal{P}_{2}(M)$ is the orthogonal complement of $T_{\rho}\mathcal{P}_{2}(M)$ with respect to the $\rho$-weighted $L^{2}$ inner product, and $P^{\rho}:L^{2}(TM,\rho)\rightarrow T_{\rho}\mathcal{P}_{2}(M)$ is the projection operator. We will use this projection operator to obtain the unique vector field defining the derivative of a map between probability spaces. The kernel of $P^{\rho}$ comprises all $\rho$-weighted divergence-free vector fields $w$ such that $\nabla\cdot(\rho w)=0$ in the distributional sense. One important property of $P^{\rho}$ is given in Proposition 3, whose complete proof is in Appendix A.
@@
2.2.4 Metric Change-of-Variables Formula
We review the following result from Riemannian geometry, which will be used to show the injectivity of the metric derivative operator (see the proof of Theorem 8). A complete proof appears in Appendix A.
@@ 2.2.4 Metric Change-of-Variables Formula
$g_{x}(\nabla(\varphi\circ f)(x),X(x))=q_{f(x)}\Big((\nabla\varphi)(f(x)),df_{x}X(x)\Big),$ (9)
where $df_{x}:T_{x}M\to T_{f(x)}N$ is the derivative operator of $f$ at position $x$.
@@ 3.1 Differentiable Curves in $\mathcal{P}_{2}(M)$
On the Lagrangian level, given an initial condition $x_{0}$, the flow $\phi_{t}(x_{0})$ generates a differentiable curve $x(t)=\phi_{t}(x_{0})$ in $M$. Similarly, on the Eulerian level, we want to have a differentiable curve $\rho_{t}$ and a vector field $v_{t}\in T_{\rho_{t}}\mathcal{P}_{2}(M)$ such that they satisfy the continuity equation (4). To begin with, we define the notion of a vector field along a curve in $\mathcal{P}_{2}(M)$.
@@
Consider a curve $\rho_{t}\in\mathcal{P}_{2}(M)$. We say that $v_{t}:[0,1]\rightarrow T_{\rho_{t}}\mathcal{P}_{2}(M)$ is a vector field along the curve $\rho_{t}$ if the tuple $(\rho_{t},v_{t})$ satisfies the continuity equation (4). If such a vector field exists, we denote it by $v_{t}:=\frac{d}{dt}\rho_{t}$.
@@
The proof appears in Appendix A and relies on the following Proposition: